OpenStudy (lukecrayonz):

Shannon who is paid time-and-a-half for hours worked in excess of 40 hours, had gross weekly wages of $391 for 44 hours worked. What is her regular hourly rate?

OpenStudy (lukecrayonz):

@dpaInc do you ever sign off? hahaha

OpenStudy (lukecrayonz):

OH, excess of 40 hours, sorry sorry

OpenStudy (anonymous):

X = hourly rate 40*X+ 4*(X*1.5) = 391 Solve for x.

OpenStudy (anonymous):

Let x be her regular hourly rate. 40x + (44-40) (1.5)x = 391 40x + 6x = 391 46x = 391 x = ..

OpenStudy (lukecrayonz):

17/2 :o, 8.5